nickblade990:

Find the measure of angle y. Round your answer to the nearest hundredth. (please type the numerical answer only) A right triangle is shown with one leg measuring 4, and the hypotenuse is labeled 20. An angle across from the leg measuring 4 is marked y degrees.

nickblade990:

i would appricate any help

Alexis1415:

Do you have the graph and/or picture that I could see?

nickblade990:

i will take a picture of it

nickblade990:

nickblade990:

thank you i really need the help

Alexis1415:

The required angle would be 11.54 degrees, Why I would say this is because, From the given triangle we have been given Angle OP is 20 units, which would be Hypotenuse. Angle AP is 4 units which would be opposite. We use sin so we would do, sin y= 4/20 sin y= 0.2 Then we would use arc-sin y=arc-sin (or just arcsin) 0.2 So, y would equal 11.54

nickblade990:

that makes sense thank you i really appricate this i hope you don't a good day, but a great day lol thx again

Alexis1415:

@nickblade990 wrote:
that makes sense thank you i really appricate this i hope you don't a good day, but a great day lol thx again
You too, and ofc. I'm glad to be of some help c:

nickblade990:

well im glad you are :)